20170502 » History » Version 6
Boyd, Steven, 02 May 2017 12:54
1 | 1 | Boyd, Steven | h1. Ops/Comm Scheduling meeting : 2/5/2017 |
---|---|---|---|
2 | |||
3 | Location : CR21 at 14:00 pm |
||
4 | |||
5 | Phone access : *New* phonebridge number |
||
6 | UK: 08082380274 |
||
7 | US: 631.621.5253 |
||
8 | Access code: 1779339517 |
||
9 | |||
10 | h2. Agenda |
||
11 | |||
12 | * Actions |
||
13 | ** Macwaters & Boehm - Arrange to install new Sumitomo compressor adsorbers - stands |
||
14 | ** Macwaters - Re-make dodgy compressor2 remote cable connector - stands |
||
15 | ** Macwaters - Prepare new F70 compressors for use. Remove old overheated CSW71 to R9. Speak to Sumitomo on April 5th regards how to repair. - in progress |
||
16 | ** Macwaters - Replace old vac sensors and new valve on cryo 1&2 - in progress |
||
17 | ** Ajit - Implement MOM/DC state machine override function for all systems; also need documentation and training for operators - documentation and training in progress |
||
18 | ** Macwaters - Report on status of FC psu repair (stands) |
||
19 | *** *This needs to be cross-checked* - e.g. Josef Boehm to check with Trevor. What is plan when it does return - will we want to test it _in situ_ before the run? |
||
20 | 6 | Boyd, Steven | *** Josef :''' "we just run 50A through it when it is connected to the magnet" ''' |
21 | 1 | Boyd, Steven | ** Hall Manager - water system change-over procedure (stands) |
22 | ** EOverton - Status of @miceraid5@ disk cleanup (stands) |
||
23 | ** JBoehm & Macwaters - get the West Wall compressors correctly labelled, as various have been re-purposed (stands) |
||
24 | ** CWhyte to contact DL to clarify their [[Hall_Task_List_Spring17|Hall Task List]] items |
||
25 | ** The Picologger has been disconnected. SBoyd to check with CWhyte whether we want it or not |
||
26 | 6 | Boyd, Steven | *** '''Contacted Colin/Josef : Josef will meet with John on Friday to discuss this''' |
27 | 1 | Boyd, Steven | ** HNebrensky to send BLOC-chasing email |
28 | |||
29 | * DC Update |
||
30 | * LH2 Update |
||
31 | 2 | Boyd, Steven | * Tracker Update |
32 | 1 | Boyd, Steven | * [[Hall_Task_List_Spring17|List of Hall activities in the shutdown]] (Colin) |
33 | * Schedule for startup |
||
34 | ** Startup run on Monday May 8th - Hall should be reserved for operations activities |
||
35 | 5 | Boyd, Steven | ** We will *not* be running the SS magnets this user cycle, or taking data with Neon. |
36 | ** We will need to run the FC at some point. |
||
37 | ** Co-ordination between DC and MOM will be key. |
||
38 | 1 | Boyd, Steven | |
39 | AOB: |
||
40 | |||
41 | * Last meeting : [[20170425|Tuesday Apr 25 2017]] |
||
42 | * Next meeting : [[20170507|Tuesday May 7 2017]] |